# VLM-Pruner: Buffering for Spatial Sparsity in an Efficient VLM Centrifugal Token Pruning Paradigm

### v3

@@ 1 Introduction
As discussed above, the excessive number of visual tokens not only reduces computational efficiency but also introduces significant redundancy: 1) Cross-modal redundancy, where only visual regions relevant to the textual query are necessary; and 2) Intra-modal redundancy, where tokens representing similar regions of the same object carry overlapping information. Recent studies have addressed these issues through training-free token pruning methods that select a representative subset of visual tokens, typically focusing on two strategies: 1) importance-based pruning [6, 3, 37, 42], which leverages intra- or cross-modal attention scores to retain salient tokens; and 2) redundancy-based pruning [35, 2], which employs greedy algorithms to iteratively choose tokens with low similarity to those selected.
@@ 2.1 Efficient Vision-Language Models
Vision–Language Models (VLMs) [9] align visual encoders with large language models to support captioning, VQA, grounding, and reasoning, with image-based [20, 18, 19, 33] and video-based [43] variants. A common recipe couples LLMs, which perform strong at text processing, with high-capacity visual encoders such as ViT [9], CLIP-ViT [28], or SigLIP [39] via cross-modal alignment.
@@ 2.2 Token Pruning
Token pruning reduces the quadratic cost of attention by compressing tokens at their source. ToMe [4] pioneered progressive token merging for ViTs [9], inspiring VLM pruning methods that are either training-required or training-free. Among training-required approaches, ATP-LLaVA [38] learns layer-wise, instance-specific pruning ratios via MLPs, but adds training overhead. We focus on training-free pruning, which falls into importance-based and redundancy-based families.
Importance-driven methods keep tokens with high (self- or cross-modal) attention: FastV [6] uses self-attention on visual tokens; SparseVLM [42] jointly prunes visual/text tokens via cross-modal attention; HiRED [3] and MustDrop [21] favor tokens strongly attending to [CLS]; PDrop [37] retains those attentive to the last-instruction token. Redundancy-reduction methods prefer diversity: DART [35] incrementally selects tokens with low similarity to text–visual anchors, while DivPrune [2] greedily preserves tokens least similar to those already kept.
@@ 4.3 Comparative Experiment
We select methods based on importance-based baselines, such as FastV [6], SparseVLM [42], PDrop [37], and MustDrop [21], as well as methods based on redundancy-based baselines, such as DART [35] and DivPrune [2], as the comprehensive baselines for comparison.
@@ 5 Conclusion
6 Additional Experiments

### v4

@@ 1 Introduction
As discussed above, the excessive number of visual tokens not only reduces computational efficiency but also introduces significant redundancy: 1) Cross-modal redundancy, where only visual regions relevant to the textual query are necessary; and 2) Intra-modal redundancy, where tokens representing similar regions of the same object carry overlapping information. Recent studies tackle these issues with training-free token pruning that selects a representative subset of visual tokens, typically focusing on two strategies: 1) importance-driven pruning [6, 3, 40, 46, 45], which leverages intra- or cross-modal attention scores to retain salient tokens; and 2) redundancy-reduction pruning [38, 2, 7, 16], which employs greedy algorithms to iteratively choose tokens with low similarity to those selected.
@@ 2.1 Efficient Vision-Language Models
VLMs [10] align visual encoders with LLMs for captioning, VQA, grounding, and reasoning, spanning image-based [23, 21, 22, 36] and video-based [47] variants. A common approach pairs an LLM with a high-capacity visual encoder such as ViT [10], CLIP-ViT [31], or SigLIP [42], and connects them through cross-modal alignment.
@@ 2.2 Token Pruning
Token pruning reduces quadratic attention cost by compressing tokens at the source. ToMe [4] introduced progressive token merging for ViTs [10], motivating VLM pruning methods that are training-required or training-free. ATP-LLaVA [41] learns layer-wise, instance-specific pruning ratios with MLPs but adds training overhead. We instead focus on training-free pruning, which mainly includes importance-driven and redundancy-reduction methods.
Importance-driven methods retain high-attention tokens. FastV [6] uses visual self-attention, SparseVLM [46] prunes visual and text tokens via cross-modal attention, and HiRED [3] and MustDrop [24] keep tokens that strongly attend to [CLS]. PDrop [40] selects tokens attending to the last instruction token, while CDPruner [45] ranks tokens by joint relevance to both text and visual tokens. Redundancy-reduction methods emphasize diversity. DART [38] incrementally selects tokens with low similarity to text and visual anchors, while DivPrune [2] and BTP [7] greedily keep tokens that are least similar to those already selected. SAINT [16] likewise favors dissimilar tokens.
@@ 4.3 Comparative Experiment
We select importance-driven baselines, such as FastV [6], SparseVLM [46], PDrop [40], MustDrop [24], and CDPruner [45], as well as redundancy-reduction baselines, such as DART [38], DivPrune [2], SAINT [16], and BTP [7], as the comprehensive baselines for comparison. We also adapt VLM-Pruner to Qwen3-VL-4B and conduct additional comparisons in the Supplementary Material.
@@ 5 Conclusion
6 More Comparative Experiment
6.1 More Baselines
we select CDPruner [45], BTP [7], and SAINT [16] as baselines for more theoretical and empirical comparisons with VLMPruner. To further demonstrate VLMPruner’s practical timeliness and applicability, we adapt it to the Qwen3-VL-4B vision-language model and conduct additional comparisons. As shown in Table 7, VLMPruner exhibits a clear advantage while retaining only 11.1% of visual tokens. Compared to other methods, (1) CDPruner [45] focuses on importance-driven pruning by selecting tokens based on joint relevance to both text and visual tokens, but it may miss critical local details. (2) BTP [7], while redundancy-reduction, disrupts fine-grained details early in the process with shallow-layer pruning, rendering later importance-driven pruning less effective. (3) SAINT [16], relying purely on redundancy-reduction pruning, may lose important local features due to its lack of spatial consideration. VLMPruner balances diversity and detail completeness by prioritizing local information through centrifugal expansion. This selects more relevant tokens while preserving fine-grained details, resulting in a more balanced selection process that excels in fine-grained visual tasks such as OCR and outperforming the second-best SAINT method by 2.15%.
[…]
Qwen3-VL-4B 62.97 84.02 2327 90.10 92.56 77.94 835 47.42 100.0%
[…]
SAINT 58.64 81.11 2025 83.13 85.87 63.75 566 41.59 87.38%
BTP 42.03 56.22 1424 61.00 76.95 55.45 431 13.27 62.06%
CDPruner 57.24 78.08 2042 86.59 86.12 63.39 486 35.18 84.31%
Ours 59.72 78.48 2129 88.51 85.23 68.37 573 42.60 89.53%
Table 7: Comparative experiments on image understanding are performed on Qwen3-VL-4B-Instruct.
6.2 More Visualizations
VLMPruner’s centrifugal expansion design is based on the assumption that visual tokens exhibit spatial locality, prioritizing the retention of important information within local neighborhoods. By balancing diversity and detail completeness through the combination of redundancy and spatial sparsity, VLMPruner ensures that nearby tokens are more likely to be selected, avoiding detail loss from dispersed coverage. As shown in Fig. 5, this reduces edge tokens’ number and improves detail completeness, leading significant improvements in tasks like fine-grained recognition. Firstly, because edge tokens exhibit extremely low similarity, DART and DivPrune select them far more frequently than VLMPruner. Moreover, VLMPruner preserves fine-grained details more effectively than DART and DivPrune, such as regions highlighted by the red circles in Fig. 5.
Figure 5: More visualizations of the actual pruning effects between baselines and VLM-Pruner. From left to right are VLM-Pruner, DivPrune, and DART. (a) The average number of edge tokens in VLM-Pruner is lower. (b) The token distribution in the VLM-Pruner model is more concentrated.
7 More Ablation Study
@@ 7.6 Ablation Study on the Pruning Layer $i$
The choice of the pruning layer is based on the following reasons: (1) According to the Align-KD [11], feature similarity between layers after the second layer exceeds 0.9, meaning the second layer already contains rich contextual information. Stage 3 compensates for unselected token information, minimizing information loss. (2) Pruning is performed on just one layer, resulting in low computational cost and time consumption. (3) It is easy and fast to adapt to different VLMs with default parameters. In contrast, methods like PDrop [40], BTP [7], and SAINT [16] require manual tuning of pruning layers and thresholds, making adaptation more complex.
7.7 Ablation Study on other base VLMs
To further validate the ”plug-and-play” claim of our method, we conduct additional ablation experiments on Qwen3-VL-4B. We select the most critical stages, Stage 1 and Stage 2, for partial benchmarking experiments. As shown in Table 15, the results confirm the strong advantages of the default parameters presented in the paper.
Benchmark Qwen3-VL-4B Query Value $\tau=0.6$ $0.7$ $0.9$ $\lambda=0.3$ $0.7$ $0.9$ Ours
[…]
MME 2327 2065 2055 2089 2096 2106 2074 2095 2096 2129
SQA 92.56 84.53 85.12 84.23 84.58 84.53 85.18 84.18 84.73 85.23
OCRBench 835 571 551 502 550 544 559 557 552 573
Table 15: Ablation Study on image understanding are performed on Qwen3-VL-4B-Instruct.